OpenStudy (roadjester):

73. The creation and study of new and mery massive elementary particles is an important part of contemporary physics. To create a particle of mass M requires an energy Mc^2. With enough energy, en exotic particle can be created by allowing a fast-moving proton to collide with a similar target particle. Consider a perfectly inelastic collision between two protons: an incident proton with mass m_p, kinetic energy K, and momentum magnitude p joins with an originally stationary target proton to form a single product particle of mass M. Not all the kinetic energy of the incoming proton is available to create the product particle because conservation of momentum requires that the system as a whole still must have some kinetic energy after the collision. Therefore, only a fraction of the energy of the incident particle is available to create a new particle. (a) Show that the energy available to create a product particle is given by \[\large Mc^2 = 2m_p c^2 \sqrt{1+{K\over 2m_p c^2}}\] This result shows that when the kinetic energy K of the incident proton is large compared with its rest energy m_pc^2, then M approaches \[(2m_pK)^{1\over 2}/c\]. Therefore, if the energy of the incoming proton is increased by a factor of 9, the mass you can create increases only by a factor of 3, not by a factor of 9 as would be expected. (b) This problem can be alleviated by using colliding beams as is the case in most modern accelerators. Here the total momentum of a pair of interacting particles can be zero. The center of mass can be at rest after the collision, so, in principle, all the initial kinetic energy can be used for particle creation. Show that \[\large Mc^2=2mc^2{(1+{K\over mc^2})}\] where K is the kinetic energy of each of the two identical colliding particles. Here, if K>>mc^2, we have M directly proportional to K as we would desire.

OpenStudy (roadjester):

@agent0smith @broken_symmetry @ybarrap @petiteme

OpenStudy (roadjester):

@wolfe8 @timo86m @TuringTest

OpenStudy (roadjester):

@ybarrap @agent0smith @broken_symmetry @wolfe8 @douglaswinslowcooper @timo86m

OpenStudy (anonymous):

Where are you allowed to start from for part (a)? If we're not using 4-vectors, then I'd start from this expression for centre-of-momentum (or centre-of-mass) energy: \[E^2_\text{CoM} = M^2c^4 = (\sum E)^2 - (\sum \mathbf{p})^2c^2\] Remember to vector-sum the momenta.

OpenStudy (anonymous):

You'll probably also need \[E = \sqrt{p^2c^2+m^2c^4} = \gamma mc^2\] and \[K = E-mc^2 = (\gamma-1)mc^2\] where \[\gamma = \sqrt{\frac{1}{1-\frac{v^2}{c^2}}}\] You can get by without using \(\gamma\) but it's there if you want to use it